I'm so confused...please help!
square root of 5x^2+9 = x
I've gotten as far as x^2+5x^2+9=0 but now I don't understand how to factor it out. The textbook confuses me.
Thanks.

Answer by stanbon(75887)

square root of 5x^2+9 = x
Square both sides to get:
5x^2+9 = x^2
4x^2+9=0
(2x+3i)(2x-3i)=0
x=(-3/2)i or x=(3/2)i
Cheers,
Stan H.